OpenStudy (lifeisadangerousgame):

How do I solve this? Evaluate for h(1/3). Show your work. h(x) = 2x2 - x

OpenStudy (anonymous):

Simply substitute in the given value (1/3) in place of x in the function. So 2x^2 - x -> 2(1/3)^2 - (1/3)

OpenStudy (lifeisadangerousgame):

ohh

OpenStudy (anonymous):

substitute 1/3 in place of x h(1/3) = 2(1/3)^2 - (1/3) = 2 (1/9) - (1/3) = (2/9) - (1/3) = (2-3)/9 = -1/3

OpenStudy (anonymous):

I think it should be -1/9 there at the end Harkirat ;)

OpenStudy (lifeisadangerousgame):

h(x) = 2x^2 - x = 2(1/3)^2 - 1/3 = 0.4356 - .33 = 0.1056 is that right?

OpenStudy (anonymous):

yeah, u r right, i made a typo....☺

OpenStudy (lifeisadangerousgame):

Oh, I didn't see your explanation Harkirat :P

OpenStudy (anonymous):

substitute 1/3 in place of x h(1/3) = 2(1/3)^2 - (1/3) = 2 (1/9) - (1/3) = (2/9) - (1/3) = (2-3)/9 = -1/9

OpenStudy (anonymous):

More or less LifeIsADangerousGame, by try and use rational expressions where you can to preserve accuracy. Every time you round it increases the error, and adds to the accumulative error in the end.

OpenStudy (lifeisadangerousgame):

Ration expressions like how?

OpenStudy (anonymous):

Like rather than using .3333, use 1/3. Learn the rules for fractions and for exponents and apply them when solving.

OpenStudy (lifeisadangerousgame):

ohh ok

OpenStudy (lifeisadangerousgame):

Just like Harkirat?

OpenStudy (anonymous):

Yes just like Harkirat. Also I'm not sure where those numbers came from, but they don't appear to be correct (your decimal numbers).

OpenStudy (lifeisadangerousgame):

Well I did 1/3 and my calculate got me .333333333 so I rounded it to .33

OpenStudy (anonymous):

I can see that but the 0.4356?

OpenStudy (lifeisadangerousgame):

and then I did .33 times 2, then .66 X .66 and then got .4356

OpenStudy (anonymous):

Ah I see. That's not quite right. x(y)^2 = x * (y^2) not (x*y)^2.

OpenStudy (lifeisadangerousgame):

ohhh ok, I see